OpenStudy (anonymous):

How many real number solutions does the equations have ? y=5x^2+2x-12 -one -two -infinitely many solutions -no solutions

OpenStudy (anonymous):

are you wanting to see a graph.... or na

OpenStudy (anonymous):

noo I just need the awnser ... .and expain ...don't know how to do it

OpenStudy (anonymous):

Oh, the answer is 'one' y=5*x^2+2*x-12 simplifies to y-5*x^2-2*x+12=0

OpenStudy (radar):

x=0, y=-12 x=1, y=-5 That is just two of many solutions.

OpenStudy (anonymous):

look at 5*x^2+2*x-12 moved these terms to the left 5*x^2-2*x+12 which makes it become y-5*x^2-2*x+12=0 thats explained but the answer is 'one'

OpenStudy (anonymous):

thanks ...blahblahblah how good are you at math ??

OpenStudy (anonymous):

actually radar is right but im pretty good